# Further Comments on Yablo’s Construction ††thanks: File: WAB, [Sch25], arXiv 2504.10370

### v4

@@ 2.2 Basic Discussion
The distinction between and i.e. between knee and foot, is conceptually important. In the case we have at a complete contradiction, at we have not yet constructed a contradiction. Thus, if we have at again an (as at this becomes an and we have to construct a contradiction for all (or not only for Otherwise, we have an escape possibility for Obviously, the contradiction need not be immediate at the important property is that ALL paths through all lead to a contradiction, and the simplest way is to have the contradiction immediately at - as in Yablo’s construction, and our saw blades.
See also the discussion of “Saw Blades” in [Sch22], Section 3.
As we work for a contradiction in the case x-, too, the simplest way to achieve this is to have a negative arrow and at again an This gives a chance to construct a contradiction at Of course, we have to construct a contradiction at too, as in the case we have an at
Of course, we may have branches originating at which all lead to contradictions in the case so (resp. at is possible, too.
But, for the simple construction, we need between and and at And this leads to the construction of contradictions for all (or as just mentioned above.
2.3 Values for y and z in Yablo Cells
We consider here what happens when we give and values In addition, we consider the variants and
Remark 2.3
False, True
@@ 2.3 Values for y and z in Yablo Cells
- (2.4)
@@
We first look only at the minimal requirements
has to be contradictory, requirement (2.1) above.
must not be classical (i.e. not or requirement (2.3) above.
The cases:
(2) (requirement (2.3))
Consider so for any so these cases are classical.
Consider For is classical, so is, too.
so
Thus, is the only non-classical case among
Consider In cases so too.
In case so
In case so
So is the only classical case among the
So the interesting cases left are
(1) (requirement (2.1))
We now check for local contradiction in the case i.e. must not be
may only be iff But if then only in case
Next, we have to look at and see if every path from leads to a contradiction, requirement (2.2) above. So neither nor must be
So and will not work.
Consider i.e.
Then in and so this does not work.
so this works.
so this works.
Consider i.e.
Then or and all cases work, except
In summary: in exactly the following cases: so they are excluded.
We summarize the results for the three requirements above:
@@
We look at these cases.
@@
Thus,
identical to
Thus,
@@ 2.4.1 Branching points
We have to branch on the path , otherwise, we have the trivial contradiction (and an escape possibility).
@@
This however, is possible only a finite number of times, e.g. then etc., see Example 1.6 (page 1.6). This constructs an infinite sequence of choices which offers an escape possibility.
@@ 2.4.2 Paths
As branching at is not different from branching before or after we assume for simplicity that we branch at Moreover, as argued above (” pipeline “), we may assume for simplicity that we always start contradictions at
Diagram 2.5
@@ 2.5 The Problem with Diamonds
If then y- and As the choices at and are independent, any branch -z, combined with any branch -z, must be conflicting, thus, given is negative, all branches on the left must be negative, likewise, all branches on the right must positive.
However, if is positive, the diamond -z, has to be contradictory, so not both branches may be negative.
A solution is to “synchronise” the choices at and which can be done e.g. by the formula
and This is a different type of formula, moreoever corresponding arrows from are missing. Even if we do not consider the paths for the diamond starting at we have and so we can make and false, and chose and which results in an escape possibility (we have to consider new diamonds starting at etc.)
This formula has an additional flaw, we want to speak about paths we chose, and not just their end points. This can be done with introducing additional points on the arrows, e.g. the arrow is replaced by now we can speak about the arrow we have given it a name.
We may also put such information in a background theory, which need no be negated.
But all this leads too far from the basically simple formalism of Yablo’s construction, so we will not discuss this any further.
@@
Thus, we conclude that the original version with 4 points and full recursion is beyond our scope, and the simplified version (see above) is nothing but the Yablo triangle.
Diagram 2.6
@@
Diagram 2.7
Diamond Essentials, Recursion on the Left, Simple Contradiction on the Right
@@ 3 Combining Cells
3.1 Overview
We follow Yablo’s construction.
(C1): We have to make contradictory.
We first try to make “directly” contradictory. That is, we will not first go to and make contradictory, nor will we make a disjunction at and then make each disjunct contradictory.
(This is not an important simplification, as more complicated constructions will finally use our simple one, too. See e.g. Diagram 1.2 (page 1.2) and its discussion.)
As Yablo did, we set and
The first step is done, is contradictory. Probabilistically, of the possibilities via are contradictory.
(C2) We have to make contradictory.
Now, must be contradictory, too. As we do not know which one is positive, we must make both and contradictory - just as a disjunction is false if every disjunct is false.
Consider, e.g., Again, we will make directly contradictory (as above for so is a conjunction, say (we might re-use here, but this will not be important) and add e.g. and to for the contradiction, so we have
But now, we have destroyed the contradiction at as for is a disjunction, and have to make ALL possibilities for contradictory for so far only the possibility is contradictory.
Probabilistically, of the possibilities at are open.
To create a contradiction for the possibility we introduce new contradictions by adding to so
Etc.
See below and Section 3.3.3. in [Sch23b].
Note that for (C1), i.e. for and for (C2), i.e. for we have to make for any disjunction all possibilities contradictory, and for any conjunction it suffices to make one conjunct contradictory, as and
Of course, following a path in the graph, conjunctions for will become disjunctions for one of the confusing aspects of reasoning.
(C1) and (C2) are antagonistic requirements. Satisfying (C1) creates new problems for (C2), and vice versa. In the limit, all requirements are satisfied, since all problems will solved in the next step, at the price of creating new problems, which will be satisfied in the next step.
Note that we don’t approximate, as long we are finite, we always have only in the infinite case we have
3.2 In More Detail
We now consider the inductive construction of the Yablo structure, see Diagram 3.1 (page 3.1) and Diagram 3.2 (page 3.2)
In (2) and (5), we have unrelated points and we may consider the model sets to be orthogonal. - We only indicate model sets briefly, without going into details (U will stand for the universe).
abbreviated
has to lead to a contradiction for
preparation,
was (part of) a full contradiction, this is now a partial contradiction, as the new possibility is added.
finish,
(branch has to be contradicted for add
has to lead to a contradiction for
preparation,
finish,
(branch has to be contradicted for add
(branch has to be contradicted for add
@@ 3.2 In More Detail
Summary:
@@
This property is “accidental”, and due to the fact that for any arrow there is also an arrow and property (2) holds for by prerequisite.
@@
Consider Diagram 3.3 (page 3.3). Here, the contradiction for is as above, but we begin a totally new contradiction for likewise for etc.
But, we see, the old construction of Diagram 3.1 (page 3.1) and Diagram 3.2 (page 3.2) is part of the more general construction of Diagram 3.3 (page 3.3), so the old construction is really minimal.
@@
Diagram Inductive Construction 1
Diagram 3.2
Diagram Inductive Construction 2
Diagram 3.3
Diagram Inductive Construction Core
@@ 4 Result
Consider structures of the type (like ’s), with one small addition: we also allow formulas of the type either as (stand alone), or as one of several elements of a conjunction where is of the type It is important that these additions also cooperate with our truth value
(This is more general than adding the constant TRUE, as it permits chaining of the new formulas, but we also add TRUE for terminal nodes.)
In addition, we assume that problems are attacked immediately: If we need a contradiction in the construction, we add it, and not a path leading to it. We also assume we use only one type of contradiction cell, and work not first with unsuitable cells, postponing the use of efficient cells. This covers also the “interdiction” of escape paths. See here e.g. the beginning of Section 2.2 (page 2.2), and Remark 2.1 (page 2.1), and Example 1.7 (page 1.7), with Diagram 1.2 (page 1.2). Without these assumptions, we have to consider many complications, which, however, do not present deeper insights.
4.1 A Suitable Injection Does the Job of Characterisation
@@
Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory iff the following conditions hold.
@@
There is an injection from YS to YS’, more precisely from the nodes of YS to the nodes of YS’, s.t., if there is a path from to in YS, then there is a path from to
For THIS injection, there is a valuation of the arrows in (the other arrows are not interesting) s.t., if form a contradiction in YS, then likewise for (Contradictions are preserved under
4.2 Proof
Fact 4.2
Yablo’s structure is the simplest structure showing that both and are contradictory.
@@ 4.2 Proof
Remark 4.3
Consider
We may chain such formulas together, like etc., which is impossible for the constant TRUE.
is preserved, if then too. Likewise, if and then
Consider Yablo adding
EG: so , this works well classically, but also for as
We may also continue beyond : or then too.
If we have mixed chains of and the last element will decide the result (trivial).
@@
Consider (without the other elements etc. in
Now consider the full YS’, and attribute values to the nodes as follows, respecting the values already given in (1):
If is minimal, then
Otherwise:
If dann
arrow
Note that these arrows have no value, i.e. we have no distinction between and
If then
: arrow
: arrow
: arrow
Note that arrows inside have a value, these values have to be preserved. The other arrows have no value, and we treat them as positive.
Moreover, we did not distinguish between minimal and not minimal this does not matter. If we want to distinguish we have to differentiate the last conjunct further.
- (3)
Remark 4.5
Asssume for simplicity that (arrows may be negative, too), likewise for etc.
We have paths etc., assume for simplicity that the only new intermediate node is respectively).
Let is not in
We set arbitrary, Thus, so and so
Of course, instead of going through the additional path might also go through etc. This changes nothing. We have neutralised the path not in
Diagram 4.1
Diagram
Diagram 4.2
Diagram

### v5

@@ 2.3 Values for y and z in Yablo Cells
Do we obtain infinite width and depth by suitable combinations of cells of the same type, see Section 2.1 (page 2.1), and [Sch25], Fact 2.4 or [Sch23b], Fact 1.4. See also Section 3 (page 3). This requires continuation at both and with constructions of value case See below (4.3) for an (unimportant) exception.
@@
We look at above requirements in the following order: (2.3), (2.1), (2.2). This is a tedious exercise and left to the reader, these are the results:
@@
We look at the results in the interesting cases.
@@
so
(Yablo) Thus,
identical to Thus,
@@
- (4.1)
- (4.2)
The simplified Saw Blade tooth is of the type with type above, see Section 7.6 in [Sch22], or Section 3 in [Sch23b]. (In this construction, will be (except for the start) again of type as we obtain this by transitivity, forced by condition (C1), see Construction 3.1 (page 3.1), and Diagram 3.1 (page 3.1), steps (2.3), (2.5), (2.6)). Both the original Saw Blade and the simplified one contain the Yablo construction.
See, e.g., Diagram 3.1, upper part, in [Sch23b]: has the role of in Yablo’s construction. is a Yablo cell. must be the head of a new Yablo cell, starting e.g. with has to be contradicted, so we need so must be the head of a new Yablo cell, so we have some etc.
- (4.4)
@@ 2.4.1 Branching points
Remark 2.4
is the origin of a new cell, so we have to branch at (Anyway, we have to branch on the path , otherwise, we have the trivial contradiction (and an escape possibility).)
@@
This however, is possible only a finite number of times, e.g. then etc., see Example 1.1 (page 1.1). This constructs an infinite sequence of choices which offers an escape possibility.
@@
Branching at on the path :
Consider the case Whatever the branching situation on is (just one branching, or multiple branchings), for the branch considered in , the path is supposed to create a contradiction. Suppose we have a branching at where i.e. it is an “OR”, then we may decide at otherwise, and not have a contradiction with the branch . Thus, we first chose a path via and then chose a contradicting path (or directly ). In Yablo’s construction, this is trivial, as is a direct link.
Definition 2.2
If , then there is forming a Yablo cell, i.e. all three components , , are negative, such that contains no branching point with for
We call this weak transitivity.
@@ 2.4.2 Paths
Note: need not be contradictory for , as is contradictory, thus, will be contradictory ( will be AND for ), so (C2) is satisfied. Thus, has the role of and we may invoke our rule that contradictions should happen immediately, and not be postponed.
@@
We might identify and , and impose this as an additional requirement for the structures we examine, along with “no postponement of contradictions”.
@@
As branching at is not different from branching before or after we assume for simplicity that we branch at Moreover, as argued above (“pipeline”), we may assume for simplicity that we always start contradictions at
@@ 2.5 The Problem with Diamonds
If then and . As the choices at and are independent, any branch , , combined with any branch , , must be conflicting, thus, given is negative, all branches on the left must be negative, likewise, all branches on the right must positive.
However, if is positive, the diamond , has to be contradictory, so not both branches may be negative, likewise for
A solution is to “synchronise” the choices at and Informally, this is possible by: “ and at choose the arrows or and at choose the arrows or ”.
However we express this, the negation is “ or choose at or choose at the arrow ”, which has nothing to do with the usual negations in Yablo structures. A way out might be to put the choices at and into a background theory which is not affected by negation.
Even if we consider arbitrary DNF formulas, we cannot speak about choices at and or arrows like Thus, within our framework, the diamond is NOT a suitable contradiction cell.
@@
Diagram 2.1
@@ 3 Combining Cells
We follow Yablo’s construction, and refer the reader to Section 1.3.1 (page 1.3.1) for an abstract discussion of this construction.
@@
The principle
If has to be the start of a Yablo cell, but is a terminal node, add an arrow
If has to be the start of a Yablo cell, exists, but not add (transitivity).
If will be the start of a new Yablo cell, if
Details
thus
will be the start of new Yablo cells.
By (C2), has to lead to a contradiction for
preparation: we use as part of the new Yablo cell, and add the new arrow as second part of the new cell, where is a new point.
finish: We add the new arrow as third part of the new cell.
Note: was (part of) a full contradiction, this is now a partial contradiction, as the new possibility is added.
By (C1),
(branch has to be contradicted for add thus
will be the start of a new Yablo cell.
By (C2), has to lead to a contradiction for
preparation: we use as part of the new Yablo cell, and add the new arrow as second part of the new cell, where is a new point.
finish: We add the new arrow as third part of the new cell.
By (C1),
(branch has to be contradicted for add Thus
Complete Yablo cell starting at
By (C1),
(branch has to be contradicted for add Thus
will be the start of a new Yablo cell.
Note:
Summary (in the limit):
Main properties:
@@
This property is “accidental”, and due to the fact that for any arrow there is also an arrow and (C2) holds for by prerequisite.
@@
We may begin a totally new contradiction for likewise for etc.
But the old construction of Diagram 3.1 (page 3.1) is part of this somewhat more general construction, so the old construction is really minimal.
@@
Remark 3.2
Above construction can be read differently: Suppose we have an (annotated) graph such that a node has truth value Then we can do above construction. E.g., has to be contradictory, (C1), so have two contradictory paths from to, say and some on the path Both paths have to lead to contradictions for (C2). Exactly one of them has to be on one of the paths (and not at the end), say this is (if both are on the paths, we would have the diamond). This way, we can reconstruct the Yablo construction within Of course, there may be additional branchings, as discussed in Section 2.4 (page 2.4). Next, we show how to neutralise them.
Definition 3.1
Consider an annotated graph a subset of its nodes, and the subgraph generated by in Consider now where in is identical to in restricted to the nodes in
Remark 3.3
above is identical up to logical equivalence for the nodes in to the following modification of replace for all and arbitrary, if and is terminal by and if is not terminal, by We call tis the isolation of in
Example: Let be as follows: and then
Consequently, in Remark 3.2 (page 3.2), the isolation is equivalent to the original Yablo structure.
@@
Diagram Inductive Construction
@@ 4 Result
4.1 Prerequisites (Limitations)
We assume that problems are attacked immediately: If we need a contradiction in the construction, we add it, and not a path leading to it. We also assume we use only one type of contradiction cell, and work not first with unsuitable cells, postponing the use of efficient cells. This covers also the “interdiction” of escape paths. See here e.g. the beginning of Section 2.2 (page 2.2), and Remark 2.1 (page 2.1). Without these assumptions, we have to consider many complications, which, however, do not present deeper insights.
Given these limitations, we show that Yablo’s structure is minimal for acyclic structures showing that and lead to contradictions, and that any unlabelled, directed graph permits a labelling showing the same iff Yablo’s original structure has a suitable injection into the graph in question.
4.2 A Suitable Injection From the Yablo Structure
@@
Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory if the following conditions hold.
@@
There is an injection from YS to YS’, more precisely from the nodes of YS to the nodes of YS’, s.t., if there is a path from to in YS, then there is a path from to in YS’, in particular, if is a direct successor of on such a path, there must be an arrow in YS’.
For this injection, we need the following condition
Let be the subgraph consisting of all nodes and the intermediate nodes on these paths.
We have to be able to construct a partial valuation in YS’, defined on the nodes in YS” (of course respecting the arrows in s.t. for any two nodes if there is an arrow in YS, then iff This need not hold for intermediate nodes. Thus, contradictions are preserved for the paths in YS’.
Definition 4.1
We extend this partial valuation in YS’ to a total valuation in YS’ as follows:
If is a node in set there is an arrow in Thus, we respect the arrows in YS’, and assures that the result is and not if
If is a node in YS”: We have already a formula describing the behaviour for all successors of in YS”. We now have to respect the other arrows in YS’, and set
there is an arrow in YS’, and not a successor of in above paths
Remark 4.1
- (1)
The second condition expresses “richness”, that we have enough independent arrows.
4.3 For the converse
Fact 4.3
Yablo’s structure is the simplest structure showing that both and are contradictory, and contained in all such structures.
@@ 4.3 For the converse
Suppose the graph has a valuation with an element
Proof
By Fact 4.3 (page 4.3), or directly by the construction Construction 3.1 (page 3.1) INSIDE we have an analogue of the Yablo structure in perhaps with paths instead of simple arrows, and perhaps with many unnecessary parts.
We can eliminate the unnecessary parts as follows, (for ease of reading with a simple example, analogously to Definition 4.1 (page 4.1)):
Suppose we have where the part is unnecessary. Then replace by so we have
Suppose we have where is unnecessary, then replace by so we have
As we have DNF formulas, this takes care of it.
5 Acknowledgements
The author would like to thank the anonymous referee for his many valuable suggestions, and the editorial board and Ms. Jane Spurr for their help.